Session: Oct 2010 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (Kaz Borkowski at Thu Oct 28 05:17:46 2010) Success:
No known problems.

Onsala (Michael Lindqvist at Thu Oct 28 05:41:56 2010) Minor failures:
Late start (around 15 UTC) due to wind.

Yebes (Pablo de Vicente at Thu Oct 28 07:07:22 2010) Minor failures: RECO FS
When Mark5B swapped between banks, at 12:40, the Mark5B hang. We rebooted the Mark5B and resumed operations. However disk pack OAN+0018 with 4 Tb has to be reset and worked with 6 of the 8 disks it has. A total of 1 hour was lost: between 12:40 and 13:30

Westerbork (Geert Kuper at Thu Oct 28 07:08:31 2010) Minor failures: PHAS
From UT 11:30 - 17:04:30 the TADU output made small jumps (10 sec based). This was caused by the fact that the receiver in RT 6 (6X), had an unstable total power output. After removing RT 6 from the adder, at 17:04:30 UT, the output of TADU is stable again. Also no data from 11:38 UT - 11:43 UT and from 17:01:00 UT - 17:04:30 UT, caused by manual backend restart.

Medicina (roda at Thu Oct 28 07:55:59 2010) Success:

Effelsberg (Uwe Bach at Thu Oct 28 12:38:24 2010) Success: